OpenStudy (anonymous):

44.99 + 29.99m(.75) Rewrite this equation by adding a 8% sales tax to the total cost Plz plz plz plz plz plz help me

OpenStudy (anonymous):

give me a minute to check myself, i think i have the explanation...

OpenStudy (anonymous):

One thing i might add i need the 8 percent to be included in the original equation

OpenStudy (skullpatrol):

What does m stand for?

OpenStudy (anonymous):

44.99 + 29.99m(.75) ok so im lost, by saying add a 8% sales cost, do you mean the entire equation? like adding 8% to the entire finished equation or to a certain part of it ?

OpenStudy (anonymous):

yea and what he said ^

OpenStudy (anonymous):

M stands for number of months

OpenStudy (anonymous):

ok....so the entire equation?

OpenStudy (jdoe0001):

of course, is months, so obvioius

OpenStudy (anonymous):

It just asked me to add the 8% to the total cost

OpenStudy (jdoe0001):

which we dunno what that is

OpenStudy (anonymous):

The original equation is what is written on the top

OpenStudy (anonymous):

Anyone here

OpenStudy (skullpatrol):

$$Total= 44.99 + 29.99m(.75) +\text{8%}[44.99 + 29.99m(.75)]$$

OpenStudy (anonymous):

Da,n that's long for seventh grade

OpenStudy (anonymous):

But I think u r right

OpenStudy (skullpatrol):

$$\Huge8\text{%}=0.08$$

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (anonymous):

Im sorry i got confused, i was preoccupied but looks like @skullpatrol got it for you